### v1

# Disentangling Task Interference within Neurons: Model Merging in Alignment with Neuronal Mechanisms

Abstract
Fine-tuning pre-trained models on targeted datasets enhances task-specific performance but often comes at the expense of generalization. Model merging techniques, which integrate multiple fine-tuned models into a single multi-task model through task arithmetic at various levels: model, layer, or parameter, offer a promising solution. However, task interference remains a fundamental challenge, leading to performance degradation and suboptimal merged models. Existing approaches largely overlook the fundamental role of individual neurons and their connectivity, resulting in a lack of interpretability in both the merging process and the merged models. In this work, we present the first study on the impact of neuronal alignment in model merging. We decompose task-specific representations into two complementary neuronal subspaces that regulate neuron sensitivity and input adaptability. Leveraging this decomposition, we introduce NeuroMerging, a novel merging framework developed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. Through extensive experiments, we demonstrate that NeuroMerging achieves superior performance compared to existing methods on multi-task benchmarks across both vision and natural language domains. Our findings highlight the importance of aligning neuronal mechanisms in model merging, offering new insights into mitigating task interference and improving knowledge fusion.
Disentangling Task Interference within Neurons: Model Merging in Alignment with Neuronal Mechanisms
Zitao Fang1, Guodong Du2, Shuyang Yu1, Yifei Guo1, Yiwei Zhang1, Jing Li2, Ho-Kin Tang2*, Sim Kuan Goh1* 1Xiamen University Malaysia, 2Harbin Institute of Technology, Shenzhen, China ait2209071@xmu.edu.my, denghaojian@hit.edu.cn, simkuangoh@gmail.com
1 Introduction
Pre-trained models (PTMs), such as foundation models and large language models (LLMs) Vaswani et al. (2017); Achiam et al. (2023); Touvron et al. (2023), have revolutionized AI by learning rich representations from large-scale datasets. These models demonstrate general capabilities while enabling effective fine-tuning for task-specific adaptation Touvron et al. (2023). PTMs have driven significant advancements across core AI domains, including natural language processing (NLP), computer vision (CV), as well as applications in medicine, law, education Bommasani et al. (2021); Moor et al. (2023); Ray (2023). Building on this success, multi-task learning (MTL) has been a paradigm for integrating task-specific abilities into a model Fifty et al. (2021), allowing generalization across multiple specialized tasks. Nonetheless, MTL requires simultaneous training on all targeted datasets, which can be costly and pose privacy concerns. Model Merging Wortsman et al. (2022); Ilharco et al. (2023); Du et al. (2024) has recently emerged as an alternative paradigm to MTL for task adaptation, enabling the training-free integration of fine-tuned models, which are increasingly being shared publicly (e.g., on Hugging Face).
Figure 1: Impacts on neuronal subspaces decomposition of T5-Large. Retaining the orthogonal subspace while removing the parallel subspace preserves near-perfect performance across all tasks. In contrast, keeping the parallel subspace while removing the orthogonal subspace leads to a significant performance drop.
Model Merging began with weight interpolation Wortsman et al. (2022) to combine the strengths of different models and has since evolved into techniques that balance competition and cooperation within shared representation and enable task editing in weight space Du et al. (2024); Ilharco et al. (2023); Ortiz-Jimenez et al. (2024). In NLP, methods such as merging task-specific language models have been explored to build and update foundation models with multi-task capabilities Raia (2021); Wan et al. (2024); Akiba et al. (2025); Wan et al. (2025). Similarly, in CV, approaches like merging Vision Transformers (ViTs) trained on different tasks or domains have been investigated to create unified models capable of handling diverse visual tasks Kim et al. (2021); Bao et al. (2022); Wang et al. (2024). In the multi-modal space, model merging has been applied to integrate models from different modalities, such as text and images, enhancing tasks like audio-visual question answering and image captioning Sung et al. (2023); Sundar et al. (2024); Dziadzio et al. (2024). These advancements underscore model merging as a promising avenue for future research.
Table 1: Model merging with different scales and granularity levels.
Method Scale Granularity Level
Fisher Merging[NeurIPS22] Fisher Matrix Parameter
RegMean[ICLR23] Inner Product Matrix Parameter
Task Arithmetic[ICLR23] Uniformed Task
Ties-Merging[NeurIPS23] Uniformed Parameter
DARE[ICML24] $1/(1-p)$ Parameter
LoraHub[COLM24] Evolver Searched Task
AdaMerging[ICLR24] Unsupervised Optimized Layer
PCB-Merging[NeurIPS24] Balancing Matrix Parameter
NeuroMerging (Ours) L1-Norm Neuron
Existing methods for model merging primarily operate at three granularities—model-level, layer-level, or parameter-level Ilharco et al. (2023); Yang et al. (2023); Du et al. (2024)—while overlooking the fundamental role of individual neurons and their connectivity Suhaimi et al. (2022); Stelzer et al. (2021), which underpins the learning process all neural networks from Perception Rosenblatt (1958) to LLMs Touvron et al. (2023). In Figure 1, we illustrate that modifying model weights along two complementary neuronal subspaces—by removing one and retaining the other—leads to distinct impacts on task performance. Notably, one subspace preserves most of the task-specific capabilities. This observation motivates us to explore model merging at the neuronal level, which could have important implications for mitigating task interference and could yield more robust merged models.
In this work, we present the first study to examine task interference at the neuronal level. Specifically, we investigate the role of neuronal alignment in model merging, illustrated in Figure 2. We begin by decomposing task-specific representations into two complementary neuronal subspaces that regulate neuron sensitivity and input adaptability. Leveraging the insights from the decomposition, we introduce NeuroMerging, a novel merging framework designed to mitigate task interference within neuronal subspaces, enabling training-free model fusion across diverse tasks. To evaluate our approach, we conduct experiments on multi-task benchmarks across both vision and natural language domains, considering various settings, including in-domain and out-of-domain generalization. Empirically, our method outperforms existing approaches. The main contributions of our paper are as follows:
  • We present the first exploration into the impact of neuronal alignment in the merging process, introducing a decomposition of task-specific representations into two complementary neuronal subspaces.
  • Based on the insights from the neuronal subspaces, we propose NeuroMerging, a new framework developed to reduce task interference in alignment with neuronal mechanism.
  • We show that NeuroMerging achieved superior performance compared to existing approaches on multi-task benchmarks in both vision and natural language processing.
2 Related Work
[…]
Figure 2: Illustration of our proposed framework. Neuronal Task Vector: Neuronal task vectors $\tau$ for the $k^{th}$ neuron are defined as the difference between the fine-tuned and pre-trained neuron for each task. Decomposition: The pre-trained $k^{th}$ neuron is decomposed into its parallel and orthogonal complementary subspaces, followed by the projection of neuronal task vectors onto these subspaces. NeuroMerging: Our proposed NeuroMerging operates within these complementary subspaces for neuronal model merging.
Multi-task learning (MTL) Fifty et al. (2021) leverages transferable knowledge to handle multiple related tasks simultaneously. Existing MTL approaches primarily rely on architectural design or optimization strategies. Architectural-based methods, such as Mixture of Experts (MoE) Shazeer et al. (2017), introduce specialized subnetworks that dynamically route inputs to task-specific experts, effectively reducing interference. However, these methods require modifying the pretrained model structure, increasing computational complexity, and limiting scalability Liu et al. (2019); Shen et al. (2024). Optimization-based approaches, on the other hand, focus on balancing task gradients or loss functions to mitigate task conflicts during training Bai et al. (2023); Kendall et al. (2018). While these methods improve convergence, they still depend on task-specific training data, which may be impractical in real-world applications due to privacy concerns or data scarcity Liang et al. (2020). In contrast, model merging offers an alternative paradigm by integrating knowledge from multiple fine-tuned models into a single unified model without requiring additional training data or architectural modifications Wortsman et al. (2022); Ilharco et al. (2023). Notwithstanding the promising findings, a key challenge in model merging is task conflict Yadav et al. (2024); Du et al. (2024), where different tasks compete for model capacity, potentially leading to suboptimal performance.
To resolve task conflicts, existing model merging methods can be categorized into three levels based on their granularity. Model-level merging combines entire model weights, typically through averaging or weighted aggregation, but often results in performance degradation due to the loss of task-specific knowledge, as seen in methods like Task Arithmetic Ilharco et al. (2023) and LoRAHub Huang et al. (2023). Layer-level merging selectively integrates layers from different models under the assumption of shared representations; for instance, AdaMerging Yang et al. (2023) adapts layer selection to better preserve task-specific information. Parameter-level merging directly manipulates individual parameters to blend knowledge from multiple models, enhancing adaptability and robustness. Techniques such as Fisher Merging Matena and Raffel (2022), RegMean Jin et al. (2023), TIES-Merging Yadav et al. (2024), DARE Le et al. (2024), and PCB-Merging Du et al. (2024) exemplify this approach. However, existing methods largely overlook the fundamental role of neuron and neuron-level interactions in task specialization. In this work, we aim to bridge this gap.
Current neural network models, from the Perceptron invented by Rosenblatt (1958) to recent massive LLMs Touvron et al. (2023), have grown significantly in scale and complexity. Nevertheless, the core principle remains unchanged: individual neurons and their connectivity still underpin the learning process Stelzer et al. (2021); Suhaimi et al. (2022). During the pre-training and fine-tuning process, neurons are not merely passive components but active elements of the network, each contributing to learning and inference Jiang et al. (2024); Islam et al. (2023). In this work, we conduct the first in-depth study on neuronal alignment in model merging, aiming to uncover its role in preserving task-specific knowledge and its potential to mitigate task conflict.
3 Methodology
In this section, we first formalize the concept of neuronal task vector for model merging and then decompose neuronal task vectors into two complementary neuronal subspaces. Subsequently, we introduce our framework, NeuroMerging, showed in Algorithm 1, which performs merging in the neuronal subspaces.
3.1 Preliminaries
In this work, we consider the model merging of a set of $T$ task specific models, $(\theta_{1},\dots,\theta_{t},\dots,\theta_{T})$, fine-tuned from a pretrained model $\theta_{0}$. With the task vector notation, each task is defined as $\tau_{t}=\theta_{t}-\theta_{0}$. The merged model is $\bar{\theta}=\theta_{0}+\phi(\tau_{1},...,\tau_{t},...,\tau_{T})$, where $\phi(\cdot)$ represents the transformation applied to each task vector $\tau_{t}$ and followed by merging. As entries of the task vector with larger magnitudes are more relevant to task-specific adaptation, only the top $r\%$ of $\mathbf{\tau}_{t}$ with the largest magnitudes are kept, while the others are set to zero Yadav et al. (2024). The masked task vector is defined as $\tau^{\text{masked}}_{t}=m_{t}\circ\tau_{t}$, where $m_{t}$ is the mask that keeps the top $r\%$ of the elements of each task vector. In the following section, we use $\tau_{t}$ to represent the masked task vector for readability.
Input: Task-specific models $\tau_{1},\tau_{2},\dots,\tau_{T}$, pretrained model $\theta_{0}$, Mask ratio $r$
Output: Merged model $\bar{\theta}$
$\tau_{t}=m_{t}\circ\tau_{t}$ // Mask task vector based on $r$
for $k\leftarrow 1$ to $K$ do
       for $t\leftarrow 1$ to $T$ do
             $\triangleright$ Create neuronal task vector.
             $\tau^{k}_{t}=\mathbf{w}^{k}_{t}-\mathbf{w}^{k}_{0}$;
             $\triangleright$ Decompose neuronal subspaces.
             $\tau^{k}_{\parallel,t}=\mathbf{P}\tau^{k}_{t}$,   $\tau^{k}_{\perp,t}=(\mathbf{I}-\mathbf{P})\tau^{k}_{t}$;
       end for
end for
$\triangleright$ Merge neuronal task vectors.
for $k\leftarrow 1$ to $K$ do
       if Validation data is available then
             Tune $\lambda_{1}$ and $\lambda_{2}$ using the validation dataset;
      else
             $\sigma_{t}=\frac{\|\mathbf{\tau}_{t}^{masked}\|_{1}}{\|\mathbf{\tau}_{t}\|_{1}}$,   $\sigma=\max(\sigma_{1},..,\sigma_{T})$;
             $\lambda_{1}=0$,   $\lambda_{2}=\frac{1}{1-\sigma}$;
       end if
      $\overline{\tau}^{k}=\lambda_{1}\psi_{\parallel}(\tau^{k}_{\parallel,1},..,\tau^{k}_{\parallel,T})+\lambda_{2}\psi_{\perp}(\tau^{k}_{\perp,1},..,\tau^{k}_{\perp,T})$;
end for
$\triangleright$ Reconstruct the merged task vector $\overline{\tau}$ by combining the $\overline{\tau}^{k}$ for each neuron.
$\bar{\theta}=\theta_{0}+\overline{\tau}$    // final merged model
return $\bar{\theta}$;
Algorithm 1 NeuroMerging
3.2 Neuronal Task Vector
Zooming in to the neuronal level of $\tau_{t}$, we define $\mathbf{w}^{k}_{t}\in\mathbb{R}^{d}$ to represent the weight vector of $k^{th}$ neuron in model $\theta_{t}$. The task-specific adaptation of this neuron, relative to the pre-trained model, is given by the neuronal task vectors:
$\mathbf{\tau}^{k}_{t}=\mathbf{w}^{k}_{t}-\mathbf{w}^{k}_{0},$ (1)
where $\mathbf{w}^{k}_{0}$ corresponds to the weight of the $k^{th}$ neuron in the pre-trained model $\theta_{0}$. Figure 2 illustrates our proposed neuronal task vector.
3.3 Neuronal Subspace Decomposition
To examine how task-specific modifications impact individual neurons, we decompose the neuronal task vectors into two complementary neuronal subspaces, visualized in Figure 2. Mathematically, this decomposition is formulated as:
$\tau^{k}_{t}=\tau^{k}_{\parallel,t}+\tau^{k}_{\perp,t},$ (2)
where $\tau^{k}_{\parallel,t}=\mathbf{P}\tau^{k}_{t}$ projects the neuronal task vectors onto the pre-trained model’s weight space, Parallel Subspace ($\mathcal{P}$), capturing neuron sensitivity. Here, $\mathbf{P}$ is the projection matrix onto the span of $\mathbf{W}^{k}_{0}$. $\tau^{k}_{\perp,t}=(\mathbf{I}-\mathbf{P})\tau^{k}_{t}$ captures the complementary orthogonal modifications, in the Orthogonal Subspace ($\mathcal{O}$), reflecting task-specific adaptability. The role of each complementary subspace:
  • Parallel Subspace ($\mathcal{P}$): this subspace captures transformations that preserve shared representations with the $\mathbf{w}^{k}_{0}$. It is also closely related to neuron sensitivity with larger magnitudes corresponding to higher sensitivity to changes in input activations.
  • Orthogonal Subspace ($\mathcal{O}$): The orthogonal complementary subspace of $\mathcal{P}$ represents novel task-specific adaptations introduced during fine-tuning, capturing input adaptability to task specific representation.
The impact of weight editing in these complementary subspace on task adaption is discussed in Section 6.3.
3.4 NeuroMerging
Our proposed NeuroMerging merges neuronal task vectors along two complementary neuronal spaces:
$\overline{\tau}^{k}$ $=\lambda_{1}\psi_{\parallel}\left(\tau^{k}_{\parallel,1},\dots,\tau^{k}_{\parallel,t},\dots,\tau^{k}_{\parallel,T}\right)$ (3)
$\quad+\lambda_{2}\psi_{\perp}\left(\tau^{k}_{\perp,1},\dots,\tau^{k}_{\perp,t},\dots,\tau^{k}_{\perp,T}\right)$
where $\psi_{\parallel}(\cdot)$ denotes the merging function for $\tau^{k}_{\parallel,t}\in\mathbb{R}$, which can be a commonly used weighted average, TIES’s disjoint merge Yadav et al. (2024), and others. $\psi_{\parallel}(\cdot)$ is also applied to non-neuronal parameters such as bias and pre-norm. For $\tau^{k}_{\perp,t}\in\mathbb{R}^{T}$, we first find dominant orthogonal subspaces within $\mathcal{O}$ using singular value decomposition (SVD) with rank being the number of tasks interact within the same neuron, and then project $\tau^{k}_{\perp,t}$ along each dimension of the SVD subspace, before applying $\psi_{\parallel}(\cdot)$ to them.
$\lambda_{1}$ and $\lambda_{2}$ are scaling parameters. When validation data is available $\lambda_{1}$ and $\lambda_{2}$ could be tuned. However, when validation is unavailable, we propose setting $\lambda_{1}=0$ (as the corresponding subspace is observed to have little impact in Section 6.4) and estimate $\lambda_{2}$ based on the impact of top $r\%$ mentioned in Section 3.1. Specifically, $\lambda_{2}$ is estimated as $\lambda_{2}=\frac{1}{1-\sigma}$, where $\sigma=max(\sigma_{1},...,\sigma_{t},...,\sigma_{T})$ and $\sigma_{t}$ is the ratio of the $L_{1}-Norm$ of the zeroed-out elements in the task vector in Section 3.1 to the $L_{1}-Norm$ of the original task vector: $\sigma_{t}=\frac{\|\mathbf{\tau}_{t}^{masked}\|_{1}}{\|\mathbf{\tau}_{t}\|_{1}}$. Detailed discussion on the parameters is provided in Section 6.4.
Subsequently, we reconstruct the task vector $\overline{\tau}$ with the all the merged neuronal task vectors $\overline{\tau}_{k}$. Finally, we obtained the final merged model $\overline{\theta}=\theta_{0}+\overline{\tau}_{rescaled}$.
4 Experimental Setup
Baseline Methods. Our baselines comprise two main categories: (1) non-model merging approaches, which include individually fine-tuned models and a multitask model trained jointly on the combined dataset serving as our theoretical upper bound, and (2) various advanced model merging techniques, including Simple Averaging (Wortsman et al., 2022), Fisher Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), Task Arithmetic (Ilharco et al., 2023), TIES-Merging-Merging (Yadav et al., 2024), PCB-Merging (Du et al., 2024), and our proposed NeuroMerging method. We report average accuracy across all tasks’ test sets as our primary evaluation metric.
Validation Set Availability. Previous works exhibit varying dependencies on a validation set. Fisher Merging inherently requires a validation set to compute the Fisher matrix. Other approaches may optionally utilize validation data for hyperparameter tuning, while RegMean leverages training data to compute and store inner product matrices for model merging. However, since these matrices match the dimensions of the original model, they introduce substantial storage and computational overhead, limiting scalability to larger models and more extensive merging tasks.
Notably, task vector-based approaches such as Task Arithmetic, Ties-Merging, and PCB-Merging, along with our proposed NeuroMerging, are substantially more lightweight and efficient. These approaches are training-free and do not rely on a validation set, making them highly practical for real-world applications. To further evaluate this advantage, we conducted additional experiments comparing task vector-based methods in scenarios where validation sets were unavailable.
Table 2: Test set performance when merging T5-Large models on seven NLP tasks.
Task($\rightarrow$) Test Set Performance
Method($\downarrow$) Validation Average paws qasc quartz story_cloze wiki_qa winogrande wsc
Zeroshot - 53.1 58.2 54.2 54.1 54.3 70.9 49.2 63.9
Finetuned - 88.0 94.4 97.1 85.3 91.0 95.7 71.6 80.6
Multitask - 88.1 94.2 98.5 89.3 92.0 95.4 73.5 73.6
Averaging[ICML22] 51.6 59.2 26.3 69.6 53.8 67.3 49.1 36.1
Task Arithmetic[ICLR23] 59.7 60.9 31.7 57.8 73.0 73.5 55.7 65.3
TIES-Merging[NeurIPS23] 76.7 80.8 92.4 77.7 81.9 78.4 61.9 63.9
PCB-Merging[NeurIPS24] 76.9 82.9 93.2 79.0 84.4 75.6 63.5 59.7
NeuroMerging (Ours) 77.6 81.1 94.3 81.6 84.7 81.2 56.7 83.9
Fisher Merging[NeurIPS22] 68.7 68.4 83.0 65.5 62.4 94.1 58.2 49.2
RegMean[ICLR23] 79.8 83.9 97.2 73.2 82.6 94.1 63.2 64.4
Task Arithmetic[ICLR23] 74.6 72.7 91.3 76.4 85.6 74.4 61.0 61.1
TIES-Merging[NeurIPS23] 79.5 82.6 94.9 72.8 87.4 85.2 66.6 66.7
PCB-Merging[NeurIPS24] 81.0 87.0 95.2 76.4 88.1 88.4 64.3 68.1
NeuroMerging (Ours) 82.2 86.4 94.3 75.9 87.9 91.2 65.9 73.6
Hyperparameters. In the absence of an additional validation set, we set $\lambda=1$ as the default value for all task-vector-based methods. For TIES-Merging and PCB-Merging, which require a masking ratio, we follow the settings of Yadav et al. (2024) and Du et al. (2024), applying $r=0.2$ as the default value across all experiments. For NeuroMerging, we set a default masking ratio of $r=0.15$, with $\lambda_{1}$ fixed at 0, while $\lambda_{2}$ is automatically adjusted according to the methodology described in Section 3.4.
When validation is allowed, we configure the non-diagonal multiplier $\alpha$ in RegMean to $0.9$, except for the T5-base model, where it is set to $0.1$. For Task Arithmetic, we perform a grid search over $\lambda$ ranging from $0.2$ to $1.5$ with a step size of $0.1$. For TIES-Merging, PCB-Merging, and NeuroMerging, we search for the optimal masking ratio r in the range $[0.05,0.2]$ with a step size of $0.05$, and $\lambda$ ($\lambda_{2}$ for NeuroMerging) from $0.8$ to $5.0$ with a step size of $0.1$.
5 Results
5.1 Merging NLP Models
Following the experimental settings from Yadav et al. (2024), we use the T5-base and T5-large models (Raffel et al., 2020), which are encoder-decoder transformers (Vaswani et al., 2017) pretrained via masked language modeling on a large text corpus, and fine-tune them independently on seven tasks: Khot et al.’s (2019) QASC, Yang et al.’s (2015) WikiQA, and Tafjord et al.’s (2019) QuaRTz for Question Answering; Zhang et al.’s (2019) PAWS for Paraphrase Identification; Sharma et al.’s (2018) Story Cloze for Sentence Completion; and Sakaguchi et al.’s (2021) Winogrande together with Levesque et al.’s (2012) WSC for Coreference Resolution. Table 2 and 7 demonstrated that our approach has superior performance over state-of-the-art methods, achieving improvements of $0.6\%$ and $1.2\%$ for T5-base and T5-large, respectively. Moreover, NeuroMerging without validation showed even more substantial gains, surpassing previous methods by $1.4\%$ for T5-base and $0.7\%$ for T5-large. For comprehensive results across all tasks and model variants, see Appendix B.2.
Figure 3: In-domain v.s. Out-domain performance.
5.2 Out-of-Domain Generalization
Building upon the experimental setup of Yadav et al. (2024), we also investigated how merged models between tasks enhance generalization in different domains. Following the approach used in prior NLP models, we merged models on seven in-domain datasets and evaluated their performance on six held-out datasets from the T0 mixture (Sanh et al., 2022) to assess cross-task generalization. These datasets encompass diverse tasks, covering three Question Answering datasets: Huang et al.’s (2019) Cosmos QA, Sap et al.’s (2019) Social IQA, and Rogers et al.’s (2020) QuAIL; one Word Sense Disambiguation dataset: Pilehvar and Camacho-Collados’s (2019) WiC; and two Sentence Completion datasets: Gordon et al.’s (2012) COPA and Zellers et al.’s (2019) H-SWAG. As shown in Figure 3, NeuroMerging outperformed the strongest baseline by $0.7\%$ and $0.5\%$ for T5-base and T5-large models, respectively, showcasing enhanced out-of-domain generalization capabilities. For more comprehensive results, please refer to Appendix B.2 Table 10 and 11.
5.3 Merging LLMs
We follow the experimental setup of Du et al. (2024) and extend our NeuroMerging to larger LLMs. Specifically, we merged three domain-specialized Llama-2-7b models (Touvron et al., 2023), and each is fine-tuned for distinct capabilities: Chinese language understanding [^1] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b, mathematical reasoning [^2] https://huggingface.co/meta-math/MetaMath-7B-V1.0 (Yu et al., 2024), and code generation [^3] https://huggingface.co/qualis2006/llama-2-7b-int4-python-code-18k (Roziere et al., 2023). To rigorously evaluate the performance of each specialized model, we employed established benchmarks tailored to their respective domains: Li et al.’s (2024) CMMLU for assessing Chinese language proficiency, Cobbe et al.’s (2021) GSM8K for mathematical capabilities, and Chen et al.’s (2021) HumanEval for code generation competency. As demonstrated in Table 3, our NeuroMerging approach exhibited substantial performance improvements, surpassing the strongest baseline by $0.9\%$. Notably, our method even outperformed the PCB-Merging utilizing Evolution Strategies (ES) optimization algorithm by $0.7\%$, underscoring the effectiveness of our proposed methodology.
Table 3: Performance comparison on LLaMA2.
Model CMMLU GSM8K Human-Eval Average
Chinese 38.6 2.3 13.4 18.1
Math 31.2 65.6 0.0 32.3
Code 33.3 0.0 17.1 16.8
Averaging[ICML22] 35.6 48.5 6.7 30.3
Task Arithmetic[ICLR23] 35.4 46.1 9.8 30.4
TIES-Merging[NeurIPS23] 36.5 53.4 12.8 34.3
PCB-Merging[NeurIPS24] 36.4 52.3 16.5 35.1
PCB-Merging+ES [NeurIPS24] 36.4 53.1 16.5 35.3
NeuroMerging (Ours) 36.1 57.2 14.6 36.0
Table 4: Performance comparison on ViT.
Method Validation ViT-B/32 Avg. ViT-L/14 Avg.
Individual - 90.5 94.2
Multi-task - 88.9 93.5
Averaging[ICML22] 65.8 79.6
Task Arithmetic[ICLR23] 60.4 83.3
TIES-Merging[NeurIPS23] 72.4 86.0
PCB-Merging[NeurIPS24] 75.9 86.9
NeuroMerging (Ours) 76.4 87.9
Fisher Merging[NeurIPS22] 68.3 82.2
RegMean[ICLR23] 71.8 83.7
Task Arithmetic[ICLR23] 70.1 84.5
TIES-Merging[NeurIPS23] 73.6 86.0
PCB-Merging[NeurIPS24] 76.3 87.5
NeuroMerging (Ours) 76.5 88.3
5.4 Merging Vision Models
We also examined the modality of vision by adhering to the experimental setup outlined by Ilharco et al. (2022, 2023). Specifically, we adopted two variants of the CLIP model (Radford et al., 2021), ViT-B/32 and ViT-L/14, as visual encoders (Dosovitskiy et al., 2021). During the fine-tuning process, we optimized the visual encoder across eight distinct tasks while maintaining a fixed text encoder configuration. This comprehensive evaluation spans multiple classification domains, encompassing remote sensing, traffic analysis, and satellite imagery recognition, with evaluations conducted on standard benchmark datasets, including Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), and SVHN (Netzer et al., 2011). Table 4 presented the results of NeuroMerging, demonstrating its competitive performance across different validation scenarios. When employing validation data, our method achieved performance improvements of $0.2\%$ for ViT-B/32 and $0.8\%$ for ViT-L/14 over state-of-the-art baselines. In the absence of additional validation, NeuroMerging further improved upon the strongest baseline by $0.5\%$ and $1.0\%$ for ViT-B/32 and ViT-L/14, respectively. These results substantiated the broad model compatibility of our approach.
6 Additional Results and Analysis
6.1 Merging without Validation Sets
When validation data is unavailable, we examine the parameters $\lambda_{1}$ and $\lambda_{2}$ selected according to Section 3.4, where $\lambda_{1}$ is set to zero, as the corresponding subspace is observed to have little impact in Section 6.4. The value of $\lambda_{2}$ is computed based on the L1-norm of masked and unmasked task vectors. Figure 7 and Table 2 and Appendix Tables 2,  8, and 9 present the evaluation of NeuroMerging on NLP and CV tasks across various model sizes, comparing it with existing methods. NeuroMerging achieves the highest average accuracy across all tasks. This demonstrates that our proposed method, with a simple rescaling, outperforms existing methods on average. Specifically, it achieved a 1.4% and 0.7% improvement over the strongest baseline for T5-Base and T5-Large, respectively. For vision models, it outperforms the strongest baseline by 0.5% and 1.0% for ViT-B/32 and ViT-L/14, respectively.
Table 5: Comparison of $\psi(\cdot)$ on Avg. Accuracy.
Method Avg. Acc
elect + mean 82.2
elect + sum 79.6
mean 79.7
sum 79.8
6.2 Ablation Study on Merging Functions
We conducted ablation experiments on various merging functions $\psi(\cdot)$ to evaluate their effectiveness in combining numerics. As shown in Table 5, among all merging functions, the elect+mean approach from TIES-Merging achieves the highest performance at $82.2\%$. In comparison, using elect+sum, averaging, and sum methods resulted in performance decreases of $2.6\%$, $2.5\%$, and $2.4\%$, respectively.
6.3 Role of Neuronal Subspaces
Figure 1 and Appendix Table 12 illustrated the impacts of neuronal subspace decomposition on T5-Large. To examine the impact of each subspace, ablation was performed separately on each subspace by retaining one while removing the other. Retaining the orthogonal subspace while removing the parallel subspace preserved near-perfect performance of finetuned checkpoints or even improved them across most tasks for T5-Large, achieving 88.0% in-domain, 53.9% out-of-domain, and an average of 58.8%. In contrast, keeping the parallel subspace while removing the orthogonal subspace resulted in a significant performance drop.
Table 6: In/Out-domain Performance of T5-large.
T5-large In-domain Out-domain Total Average
Fine-tuned 88.0 53.8 58.7
Keep Orthogonal 88.0 53.9 58.8
Keep Parallel 52.7 51.8 51.9
6.4 Robustness of Hyperparameters
We systematically investigated the impact of hyperparameters on merging performance: $\lambda_{1}$ and $\lambda_{2}$, which control the parallel and orthogonal subspace contributions, respectively, and the mask ratio $r$.
Relationship Between $\lambda_{1}$ and $\lambda_{2}.$ To examine the effects of $\lambda_{1}$ and $\lambda_{2}$, we conducted a grid search with $\lambda_{1}\in[0,1.0]$ and $\lambda_{2}\in[3.0,4.0]$ at $0.1$ intervals, fixing $r=10\%$. As visualized in Figure 4, the performance exhibited column-wise uniformity in the heatmap, indicating insensitivity to variations in $\lambda_{1}$, which aligns with our earlier discussion on the role of the orthogonal subspace in Section 6.3. The optimal performance occurred at $\lambda_{2}=3.6$, attributed to the substantial proportion of masked variables.
Figure 4: Impacts of $\lambda_{1}$ and $\lambda_{2}$ on T5-Large.
Masking Ratio $r$. We observed robust performance across different ratios, with accuracy peaks at $15\%$ and $10\%$ for T5-Base and T5-Large, respectively. Performance variations remain bounded (within $2.5\%$ for T5-Base and $4\%$ for T5-Large) and stabilize beyond $35\%$.
7 Conclusion
In this paper, we revisited model merging from the core principle neuron and its connectivity that underpin the learning process of recent deep neural networks and LLMs. Specifically, we presented the first exploration into the impact of neuronal alignment in the model merging process, by decomposing of task-specific representations into two complementary neuronal subspaces and examining their roles. Based on these insights, we proposed NeuroMerging, a novel framework designed to reduce task interference within neurons. Our empirical evaluations demonstrate that NeuroMerging achieves superior performance compared to existing approaches on multi-task benchmarks across both vision and natural language processing tasks.
Future work could extend NeuroMerging to larger models or multimodal architectures with more tasks. While our study introduces a neuronal mechanistic perspective on task interference in model merging, it captures only one aspect of neuronal mechanisms, as it does not yet explore interactions between synchronized neuron groups or neuronal dynamics. Further research is needed to investigate these factors and deepen our understanding of neuronal mechanisms in model merging or even multi-task learning.
8 Limitations
While our work provides the first neuronal mechanistic perspective on task interference when merging language models, (1) it remains a partial view of neuronal mechanisms as it does not yet explore interactions between synchronized neuron groups or neuronal dynamics during merging and inference, which require further investigation. Moreover, this work shares similar limitations with current SOTA model merging methods, including (2) the effectiveness of task arithmetic in model merging relies on selecting fine-tuned checkpoints that are beneficial for specific domains, ensuring they originate from the same pretrained model, and addressing hyperparameter sensitivity; and (3) More effort is needed to develop a mathematical understanding of why and when task arithmetic in model merging works well, despite its simplicity and efficiency.
Appendix A Statistical Analysis of Parameters
From the statistical analysis of neuronal versus non-neuronal parameters, it is observed that neuronal parameters dominate the T5-Base and T5-Large, showed in Figure. 5. As a result, they play a major role in shaping the model’s learning dynamics, task adaptability, and overall performance. This dominance suggests that understanding and optimizing neuronal parameter interactions is crucial for improving model merging, reducing task interference, and enhancing generalization across diverse tasks.
Figure 5: Statistical analysis of neuronal versus non-neuronal parameters.
Appendix B Additional Results
B.1 Mask Ratio vs. Performance
When validation is available, we analyze the impact of the mask ratio on performance, showed in Figure 6. Maintaining $r$ at 10–20% improves performance, whereas exceeding this range often leads to degradation. This suggests that an optimal masking ratio balances information retention and redundancy reduction.
Figure 6: Impacts on mask ratio $r$.
B.2 Comprehensive Task-Level Results
We provide all task-level results in T5-Base, T5-Large (Raffel et al., 2020), LLaMA2 (Touvron et al., 2023), ViT-B/32, and ViT-L/14 (Dosovitskiy et al., 2021), respectively. The task-level results of the in-domain experiments for all models can be found in Tables 789. The task-level results of the out-domain experiments for T5-Base and T5-Large can be found in Tables 10 and 11. Lastly, Table 12 shows the task-level results from Section 6.3 when only one of the neuronal subspace is retained.
Table 7: Test set performance when merging T5-Base models on seven NLP tasks. Please refer to Section 5.1 for experimental details.
Task($\rightarrow$) Test Set Performance
Method($\downarrow$) Validation Average paws qasc quartz story_cloze wiki_qa winogrande wsc
Zeroshot - 53.5 49.9 35.8 53.3 48.1 76.2 50.0 61.1
Finetuned - 79.6 93.9 98.0 81.4 82.5 95.4 51.9 54.2
Multitask - 83.6 94.0 97.9 82.5 86.7 95.0 64.1 65.3
Averaging[ICML22] 64.2 65.1 81.1 59.9 48.8 94.7 50.9 48.6
Task Arithmetic[ICLR23] 60.6 78.5 30.0 56.8 65.6 95.0 49.6 48.6
TIES-Merging[NeurIPS23] 73.6 82.4 94.1 71.9 66.0 91.2 51.5 58.3
PCB-Merging[NeurIPS24] 73.4 83.1 92.6 72.6 73.4 88.0 51.5 52.8
NeuroMerging (Ours) 74.8 81.8 96.5 69.3 67.9 94.8 51.0 62.5
Fisher Merging[NeurIPS22] 68.3 66.7 85.6 63.5 57.1 90.1 54.2 60.8
RegMean[ICLR23] 72.7 77.2 93.8 63.6 64.6 90.4 58.4 60.7
Task Arithmetic[ICLR23] 73.6 83.2 89.9 69.3 72.9 95.2 52.1 52.8
TIES-Merging[NeurIPS23] 76.4 88.6 94.1 74.5 75.6 92.1 53.2 56.9
PCB-Merging[NeurIPS24] 76.9 88.2 95.2 71.0 77.3 95.1 51.9 59.7
NeuroMerging (Ours) 77.5 87.5 95.7 68.2 76.8 94.5 51.8 68.1
Table 8: Test set performance when merging ViT-B/32 models on eight vision tasks. Please refer to Section 5.4 for experimental details.
Task($\rightarrow$) Test Set Performance
Method($\downarrow$) Validation Average SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Individual - 90.5 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4
Multitask - 88.9 74.4 77.9 98.2 98.9 99.5 93.9 72.9 95.8
Averaging[ICML22] 65.8 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1
Task Arithmetic[ICLR23] 60.4 36.7 41.0 53.8 64.4 80.6 66.0 98.1 42.5
TIES-Merging[NeurIPS23] 72.4 59.8 58.6 70.7 79.7 86.2 72.1 98.3 54.2
PCB-Merging[NeurIPS24] 75.9 65.8 64.4 78.1 81.1 84.9 77.1 98.0 58.4
NeuroMerging (Ours) 76.4 64.7 64.2 77.0 83.9 86.2 78.0 98.5 58.7
Fisher Merging[NeurIPS22] 68.3 68.6 69.2 70.7 66.4 72.9 51.1 87.9 59.9
RegMean[ICLR23] 71.8 65.3 63.5 75.6 78.6 78.1 67.4 93.7 52.0
Task Arithmetic[ICLR23] 70.1 63.8 62.1 72.0 77.6 74.4 65.1 94.0 52.2
TIES-Merging[NeurIPS23] 73.6 64.8 62.9 74.3 78.9 83.1 71.4 97.6 56.2
PCB-Merging[NeurIPS24] 76.3 66.7 65.5 78.5 79.3 86.4 77.1 98.2 59.1
NeuroMerging (Ours) 76.5 65.3 65.7 77.1 84.8 84.5 77.9 98.3 58.5
Table 9: Test set performance when merging ViT-L/14 models on eight vision tasks. Please refer to Section 5.4 for experimental details.
Task($\rightarrow$) Test Set Performance
Method($\downarrow$) Validation Average SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Individual - 94.2 82.3 92.4 97.4 100 98.1 99.2 99.7 84.1
Multitask - 93.5 90.6 84.4 99.2 99.1 99.6 96.3 80.8 97.6
Averaging[ICML22] 79.6 72.1 81.6 82.6 91.9 78.2 70.7 97.1 62.8
Task Arithmetic[ICLR23] 83.3 72.5 79.2 84.5 90.6 89.2 86.5 99.1 64.3
TIES-Merging[NeurIPS23] 86.0 76.5 85.0 89.3 95.7 90.3 83.3 99.0 68.8
PCB-Merging[NeurIPS24] 86.9 75.8 86.0 89.2 96.0 88.0 90.9 99.1 70.0
NeuroMerging (Ours) 87.9 77.0 86.9 90.3 96.3 89.9 92.1 99.2 71.8
Fisher Merging[NeurIPS22] 82.2 69.2 88.6 87.5 93.5 80.6 74.8 93.3 70.0
RegMean[ICLR23] 83.7 73.3 81.8 86.1 97.0 88.0 84.2 98.5 60.8
Task Arithmetic[ICLR23] 84.5 74.1 82.1 86.7 93.8 87.9 86.8 98.9 65.6
TIES-Merging[NeurIPS23] 86.0 76.5 85.0 89.4 95.9 90.3 83.3 99.0 68.8
PCB-Merging[NeurIPS24] 87.5 76.8 86.2 89.4 96.5 88.3 91.0 98.6 73.6
NeuroMerging (Ours) 88.3 77.3 87.1 90.1 96.1 91.0 92.2 99.4 73.0
Table 10: Out-of-Distribution performance of T5-Base models checkpoints on six tasks. Please refer to Section 5.2 for experimental details.
Method ($\downarrow$) Average cosmos_qa social_iqa quail wic copa h-swag
paws 37.2 25.0 37.0 29.9 49.5 57.4 24.5
qasc 36.5 21.2 37.4 29.5 49.8 54.4 26.7
quartz 36.9 24.7 36.9 28.8 48.5 57.4 25.0
story_cloze 36.8 21.9 36.4 25.7 53.6 57.4 26.1
wiki_qa 36.2 25.9 36.3 29.9 51.2 48.5 25.2
winogrande 36.8 23.9 37.9 24.1 51.8 58.8 24.4
wsc 39.5 26.9 38.1 29.5 55.4 61.8 25.2
Pretrained 36.8 22.9 36.4 29.9 50.8 55.9 24.8
Averaging[ICML22] 37.4 23.7 36.8 29.3 51.3 58.8 24.8
Fisher Merging[NeurIPS22] 33.8 15.6 21.9 24.9 65.6 53.1 21.9
Task Arithmetic[ICLR23] 36.9 19.0 35.6 29.5 54.0 55.9 27.7
RegMean[ICLR23] 34.3 23.1 28.1 24.9 48.4 62.5 18.8
TIES-Merging[NeurIPS23] 38.5 21.9 37.4 29.3 52.0 64.7 25.5
PCB-Merging[NeurIPS24] 38.5 22.8 37.5 29.1 51.3 63.2 27.0
NeuroMerging (Ours) 39.2 21.2 37.3 29.9 52.0 69.1 25.4
Table 11: Out-of-Distribution performance of T5-Large models checkpoints on six tasks. Please refer to Section 5.2 for experimental details.
Method ($\downarrow$) Average cosmos_qa social_iqa quail wic copa h-swag
paws 38.2 28.4 37.6 25.4 60.9 51.5 25.2
qasc 37.9 23.1 37.0 25.5 49.0 64.7 28.1
quartz 36.2 26.1 38.0 25.7 51.3 50.0 26.2
story_cloze 37.9 22.9 37.5 24.5 51.2 64.7 26.6
wiki_qa 35.0 23.2 37.4 26.1 51.2 47.1 25.1
winogrande 36.1 25.2 39.6 24.1 51.3 50.0 26.4
wsc 37.2 26.2 38.8 28.8 55.4 48.5 25.8
Pretrained 36.3 23.7 37.8 28.1 51.2 51.5 25.5
Averaging[ICML22] 36.7 25.3 37.0 23.4 51.5 57.4 25.9
Fisher Merging[NeurIPS22] 32.0 34.4 25.0 26.1 40.6 56.2 9.4
Task Arithmetic[ICLR23] 39.2 24.6 38.0 27.3 58.6 58.8 28.1
RegMean[ICLR23] 36.0 34.4 28.1 25.3 62.5 50.0 15.6
TIES-Merging[NeurIPS23] 40.1 25.1 40.8 23.0 56.3 67.6 27.6
PCB-Merging[NeurIPS24] 40.4 25.6 40.7 25.7 55.1 66.2 29.3
NeuroMerging (Ours) 40.9 24.7 40.7 26.6 56.4 69.1 27.9
Table 12: Test set performance comparison of T5-Large models under different keeping strategies (naive finetuned, keep orthogonal, and keep parallel) across seven NLP tasks. Please refer to Section 1 and 6.3 for experimental details.
Task($\rightarrow$) Test Set Performance
Method($\downarrow$) Dataset ($\downarrow$) paws qasc quartz story_cloze wiki_qa winogrande wsc Average
Fine-tuned paws 94.4 18.0 53.3 53.1 87.9 49.8 54.2 58.7
qasc 54.3 97.1 55.7 64.7 66.3 49.8 41.7 61.4
quartz 59.9 65.1 85.3 51.2 72.5 48.9 62.5 63.6
story_cloze 53.4 31.9 54.2 91.0 57.4 49.1 56.9 56.3
wiki_qa 55.8 16.3 50.9 53.7 95.7 48.7 63.9 55.0
winogrande 55.7 50.2 62.4 55.3 78.4 71.6 56.9 61.5
wsc 55.8 16.3 57.7 48.5 73.3 47.4 80.6 54.2
Total Avg. In-Domain: 88.0 Out-Domain: 53.8 All: 58.7
Orthogonal paws 94.4 18.1 53.3 53.3 88.1 49.6 56.9 59.1
qasc 54.4 97.1 55.6 64.8 66.3 50.0 43.1 61.6
quartz 60.0 65.0 85.3 51.5 72.5 48.5 63.9 63.8
story_cloze 53.3 31.8 53.8 90.9 57.7 49.0 56.9 56.2
wiki_qa 55.8 16.3 51.1 53.5 95.7 48.2 63.9 54.9
winogrande 55.7 49.8 62.2 55.8 78.4 71.7 56.9 61.5
wsc 55.8 16.6 57.4 48.2 73.4 47.4 80.6 54.2
Total Avg. In-Domain: 88.0 Out-Domain: 53.9 All: 58.8
Parallel paws 54.4 14.5 53.4 54.2 71.8 49.5 63.9 51.7
qasc 55.3 14.9 54.0 54.3 71.0 48.6 63.9 51.7
quartz 55.3 14.3 55.6 54.2 71.7 49.4 63.9 52.1
story_cloze 55.3 14.3 54.1 53.8 71.1 49.3 63.9 51.7
wiki_qa 55.6 14.7 54.5 54.3 77.1 49.1 63.9 52.7
winogrande 55.3 14.8 54.2 53.6 71.8 49.5 63.9 51.9
wsc 55.3 14.7 53.7 53.7 72.5 49.5 63.9 51.9
Total Avg. In-Domain: 52.7 Out-Domain: 51.8 All: 51.9
Figure 7: Comparison of merging methods on NLP with T5-Large (Left) and CV with ViT-L/14 (Right) without validation datasets. NeuroMerging outperforms existing methods in most tasks. Please refer to Section 6.1 for more discussion.
Appendix C Implementation Details
We executed all our experiments on Nvidia A6000 GPUs equipped with 48GB RAM. The merging experiments demonstrated highly computational efficiency, with evaluation times under 2 minutes for T5-Base, T5-Large, ViT-B/32, and ViT-L/14 models. For large language model, specifically LLaMA2, the validation process across three datasets required approximately 40 minutes per complete evaluation cycle.
Appendix D Dataset Details
We utilized several datasets, and each of them comes with specific licenses. The following datasets are available under the Creative Commons License: WiC (Pilehvar and Camacho-Collados, 2019), WSC (Levesque et al., 2012), Story Cloze (Sharma et al., 2018), QuaRTz (Tafjord et al., 2019), Cars (Krause et al., 2013), and GTSRB (Stallkamp et al., 2011). Winogrande (Sakaguchi et al., 2021) and QASC (Khot et al., 2019) are distributed under the Apache License, while COPA (Gordon et al., 2012) is covered by the BSD-2 Clause License. WikiQA (Yang et al., 2015) is governed by the Microsoft Research Data License Agreement. Cosmos QA (Huang et al., 2019) is licensed under the CC BY 4.0. QuAIL (Rogers et al., 2020) and CMMLU (Li et al., 2024) are licensed under the CC BY-NC-SA 4.0. H-SWAG (Zellers et al., 2019), GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and EuroSAT (Helber et al., 2019) fall under the MIT License, and MNIST (LeCun, 1998) is licensed under the GNU General Public License.
For the datasets DTD (Cimpoi et al., 2014), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), SVHN (Netzer et al., 2011), Social IQA (Sap et al., 2019), and PAWS (Zhang et al., 2019), we were unable to determine specific licenses. However, they are publicly shared for research and education purposes.